OpenStudy (anonymous):

Change the given equation into slope-intercept form. 5x – 2y = -10

OpenStudy (amistre64):

just gotta solve for y all by its little lonesome

OpenStudy (amistre64):

or, determine the slope and the y intercept and fill them int to the y = mx+b format

OpenStudy (amistre64):

given the equation; ax + by = c; slope = -a/b , and the y int = c/b

OpenStudy (anonymous):

ya, basically or find intercepts asa: y int. : a(0) + by = c x int : ax + b(0)= c